### v2

# On the Robustness of RSMA to Adversarial BD-RIS-Induced Interference

Abstract
This article investigates the robustness of RSMA (RSMA) in multi-user MISO (MISO) systems to interference attacks against channel acquisition induced by BD-RIS. Two primary attack strategies, random and aligned interference, are proposed for fully connected and group-connected RIS (RIS) architectures. Valid random reflection coefficients are generated exploiting the Takagi factorization, while potent aligned interference attacks are achieved through optimization strategies based on a QCQP (QCQP) reformulation followed by projections onto the unitary manifold. Our numerical findings reveal that, when perfect CSI (CSI) is available, RSMA behaves similarly to SDMA (SDMA) and thus is highly susceptible to the attack, with BD-RIS inducing severe performance loss and significantly outperforming diagonal RIS. However, under imperfect CSI, RSMA consistently demonstrates significantly greater robustness than SDMA, particularly as the system’s transmit power increases.
I Introduction
[…]
Figure 1: An adversarial entity configures a BD-RIS to launch interference attacks on multiple RSMA users during downlink data transmission.
RSMA is a split transmission scheme whereby the BS (BS) encodes a part of the users’ data into private messages - transmitted via private precoders - with the remaining users’ data being encoded into a common message - transmitted via a common precoder [10, 11]. On the receiver side, SIC (SIC) is utilized to aid in decoding the individual users’ messages. The private precoders are designed as unicast precoders, which - similarly to SDMA - makes them sensitive to imperfect CSI. The common precoder is constructed as a multicast precoder and, thus, should be more robust to inter-user interference arising due to CSI imperfections. Consequently, by adaptively re-allocating the power between the private and common precoders, RSMA is able to adapt to CSI imperfection [11]. In the conference version of this article [6], we observed that RSMA is able to adapt also in the face of a channel estimation attack induced by a passive RIS, when the CSI available at the BS is imperfect. However, the extent of RSMA’s susceptibility to RIS-induced attacks is still an open challenge with many different parts of the RSMA protocol, including the uplink pilot training, power allocation, and SIC imperfections, requiring deeper investigation. One particularly poignant point is the susceptibility to attacks by novel RIS architectures that feature non-diagonal reflection matrices. These architectures, referred to as BD-RIS, generalize conventional RIS designs by introducing inter-element connections [12]. Such connections enable more advanced optimization capabilities and enhanced energy radiation, which in turn boosts reflection performance at the cost of additional circuit complexity. Nevertheless, the enhanced reflection efficiency of BD-RIS may also strengthen their adversarial capabilities, potentially leading to attacks with increased degradation impact. Identifying the distinct vulnerabilities that BD-RIS may introduce to RSMA systems and the advantages that would motivate an attacker to leverage BD-RIS constitutes a major research gap.
This article reports on our extensive study and analysis of the robustness of the RSMA protocol to attacks against channel acquisition induced by BD-RIS, under different attack modes during the uplink pilot training, varying numbers of reflecting elements, and various levels of CSI and SIC imperfections. In our study, we consider three BD-RIS element network architectures: fully connected, group-connected, and single-connected (corresponding to the diagonal RIS). For each architecture, we propose two different BD-RIS-induced interference attacks that exploit the training protocol employed at the BS, namely random and aligned. In both attacks, during uplink pilot training, the BD-RIS is configured to absorb or reflect with random reflection coefficients. Then, during data transmission, the BD-RIS is configured with newly generated random reflection coefficients (random attack) or with optimized coefficients that maximize the reflected power based on the knowledge of the RIS channels (aligned attack). In each case, the proposed algorithm finds a set of reflection coefficients that satisfy the symmetry and unitary constraints of the BD-RIS.
Our numerical study reveals that, when perfect CSI is available, RSMA behaves similarly to SDMA and thus is highly susceptible to the attack, with BD-RIS group-connected and fully connected architectures inducing larger performance degradations than a single-connected RIS. However, RSMA displays a surprising degree of robustness to the BD-RIS-induced attack when the CSI is imperfect, similarly to what we previously reported for conventional RIS in [6]. When the system is under the BD-RIS-induced attack, the adaptation mechanism of RSMA reacts to the inaccurate CSI by allocating more power to the common message. When the transmit power is low (below 10 dBm) RSMA’s rate is being severely degraded by the attack (on par with SDMA), but as the transmit power increases, RSMA is able to recover from the attack. This trend can be observed across all types of BD-RIS attacks. In order to directly quantify this inherent robustness, we propose the robustness index that we define and discuss in subsequent sections. As we gradually adjust the transmit power, the robustness index for SDMA under attack steadily decreases towards a saturation point. In contrast, RSMA’s robustness index initially declines to a local minimum before beginning a continuous rise, consistently remaining above SDMA’s robustness. This highlights the ability of RSMA to sustain performance under attack, given a sufficiently large transmit power. We also show that the performance degradation induced by group and fully connected architectures becomes increasingly severe compared to a conventional RIS as the number of reflecting elements grows. Our key contributions are as follows:
  • We propose and analyze two distinct attack strategies, employing either random or optimized reflection coefficients, for each BD-RIS architecture. These strategies are designed to rigorously satisfy the inherent constraints of the respective BD-RIS types, and are targeted at disrupting multi-user data transmission in an RSMA system.
  • For the random interference attack strategy, we leverage matrix theory to propose a simple algorithm based on Takagi factorization. This algorithm generates valid random BD-RIS reflection coefficient matrices that are symmetric and unitary (or block-wise symmetric and unitary for group-connected RIS architectures).
  • For the aligned interference attack strategy, we develop an algorithm to maximize the unwanted reflected power towards legitimate users. To address the resulting complex optimization problems, we propose to exploit the symmetric structure of the BD-RIS reflection matrices by employing duplication matrices. This approach reformulates the problems into simplified QCQP versions that are efficiently solvable using SVD (SVD), followed by projections onto the unitary manifold via the Takagi factorization.
  • We propose and investigate two metrics to quantify the robustness of RSMA systems against BD-RIS-induced attacks: rate degradation and robustness index, ranging from 0 (minimum robustness) to $1$ (maximum robustness). Our findings demonstrate that under imperfect CSI, RSMA exhibits substantially greater robustness compared to non-adaptive protocols such as SDMA. Moreover, the results highlight the scalability of RSMA, as its degradation remains consistently lower than that of SDMA, with the gap widening as the RIS dimension increases.
  • Furthermore, we reveal that BD-RIS offers a significant adversarial advantage over conventional RIS across all investigated scenarios, both under random and aligned interference attacks, an advantage that becomes particularly pronounced when the attacking RIS operates in a reflective mode during the uplink pilot training phase. At the same time, we uncover that while RSMA is fundamentally more robust to different sources of interference, its robustness strongly depends on the quality of SIC. Even small imperfections in SIC can cause a significant decrease in the system sum rate.
Notation: Boldface lower-case letters denote vectors and upper-case represent matrices. The $\ell_{2}$ norm of a vector $\mathbf{a}$ is denoted by $\|\mathbf{a}\|_{2}$, and the Frobenius norm of a matrix $\mathbf{A}$ by $\|\mathbf{A}\|_{F}$. The $i$th column of a matrix $\mathbf{A}$ is denoted by $\mathbf{A}]_{:,i}$, the transpose and Hermitian transpose of $\mathbf{A}$ are represented by $\mathbf{A}^{T}$ and $\mathbf{A}^{H}$, respectively, and $\otimes$ represents the Kronecker product. The operator $\mathrm{vec}(\cdot)$ transforms a matrix into a column vector by stacking its columns sequentially, $\mathrm{unvec}(\cdot)$ reshapes a vectorized matrix into its original dimensions, and $\mathrm{vech}(\cdot)$ is the half-vectorization operator, which stacks the elements from the lower triangular part (including the main diagonal) of a square matrix into a column vector. Moreover, $\mathrm{vecd}(\cdot)$ converts the diagonal elements of a square matrix into a column vector, $\mathrm{diag}(\cdot)$ transforms a vector into a diagonal matrix, and $\mathrm{bdiag}(\cdot)$ constructs a block diagonal matrix sequentially formed by the input matrices.
II Related Work
The concept of an adversary that configures RIS to launch attacks against the wireless link was originally proposed in [1]. The proposed attack relied on using the RIS to create cancellation signals. Although the attack has the potential to cause significant power degradation in the receiver, the attack requires highly accurate CSI, which would limit its feasibility. Notably, many works that followed showed that a RIS-induced attack can be made effective with limited or even no CSI [3, 5, 13, 6, 8], as long as it takes advantage of the fact that for many wireless protocols the CSI acquisition is sufficiently separated in time from data transmission. For instance, in [3], it was shown that the channel equalization can be disrupted by the operation of the adversarial RIS that randomly flips its reflection pattern between channel acquisition and symbol transmission. This random flipping is particularly harmful to multiple access methods that rely on multi-antenna systems, such as SDMA. In this case, adversarial RIS increases CSI inaccuracy, making linear precoders ineffective in removing inter-user interference. Exactly this kind of attack with randomly set RIS coefficients, humorously referred to by their authors as the “disco-ball attack”, was proposed in [5, 13]. The optimized version of this attack that aligns the RIS channels to boost interference was proposed and studied in [6, 8]. Other protocols and services, such as beam management [14], physical layer key generation [15, 16], and integrated sensing and communication [17, 18], have also been shown to be vulnerable to the attack with RIS flipping its coefficients between different phases of the wireless protocol. However, many open questions remain in studying RIS-induced attacks, particularly as they relate to identifying protocols and methods that may be robust or resilient to the proposed attacks.
RSMA was introduced as a multiple access strategy that exploits a split transmission mechanism to adaptively treat inter-user interference at either the BS or receiver side, depending on the accuracy of the available CSI [10, 11]. When CSI is perfect, RSMA resorts to linear precoding at the BS that can effectively take care of the inter-user interference. On the other hand, when CSI is inaccurate, RSMA employs multicast transmission and SIC at the receiver side to decode (part of) inter-user interference. There are a variety of benefits that come from this increased flexibility, such as increased spectral and energy efficiency [19], improved outage [20], and reduced latency [21, 22]. RSMA has also been shown to be robust and resilient to user mobility [23], and binary link failures [24]. However, one of the properties gaining the most attention is its inherent robustness and resilience to inaccurate CSI [11], which can also be quantified, following [24]. Since RIS-induced interference attacks inherently rely on manipulating the CSI, the outstanding question is how robust RSMA is to inaccurate CSI that results from adversarial RIS activities.
In the preliminary work described in the conference version of this article, we considered the impact of adversarial diagonal RIS on RSMA [9]. What we observed is the surprising robustness of RSMA to the RIS-induced attack, which manifests itself under imperfect CSI scenarios. In this article, we extend our preliminary study by directly quantifying the robustness of RSMA to adversarial attacks using a new metric that we refer to as the robustness index. We focus our attention on the impact of an attack with a BD-RIS architecture that generalizes conventional (diagonal) RIS. Alternative RIS architectures were considered for adversarial attacks to create non-reciprocal channels [25], and increase their angular span [26]. However, this is the first work that considers BD-RIS and inter-element connections in the context of an adversarial scenario. Consequently, in this paper, we propose a set of new algorithms that find valid reflection coefficients that obey the properties of BD-RIS. Furthermore, we consider different attack modes during the uplink pilot training, and the impact of adaptive power allocation, and SIC imperfections.
Figure 2: Simplified diagram of the RIS-induced attack strategy.
III System Model
Consider the multi-user MISO system depicted in Fig. 1, where a BS equipped with a linear array of $M$ antennas communicates with $U$ single-antenna users, represented by the index set $\mathcal{U}=\{1,2,\cdots,U\}$, with the aid of RSMA. The system operates in the presence of an adversarial entity that controls a BD-RIS comprising $D$ reflecting elements. The attacker manipulates the BD-RIS to disrupt the channel estimation phase at the BS, configuring it to either absorb incoming pilots [6, 5] (absorption mode) or reflect them randomly [13] (reflective mode). The impact of the mode choice will be investigated in Section V.
Subsequently, the attacker generates a new set of reflecting coefficients during the data transmission phase, altering the propagation environment in a manner unknown to the BS. This ensures that the BS operates with corrupted CSI, as its estimates from the training phase do not reflect the RIS’s true, active configuration during data transmission. The discrepancy between estimated and actual CSI reduces the performance of precoders, ultimately generating inter-user interference. Fig. 2 illustrates the overall adversarial protocol.
Note that the attacker must ensure that its RIS configurations differ between the uplink training and downlink data transmission phases. As explained in [13], this can be accomplished by generating a new set of coefficients on a timescale comparable to or shorter than the pilot training duration. A more sophisticated alternative is for the adversary to actively synchronize its attack. Since pilot sequences and their transmission intervals are typically public information defined by the wireless standard (e.g., in LTE Advanced and 5G NR (NR) systems), an attacker can passively monitor the channel to detect these periodic transmissions and determine the timing of the uplink training phase [27].
III-A BD-RIS Architectures
We consider two BD-RIS types, the fully connected and group-connected architectures. In the fully connected RIS, all elements are interconnected through reconfigurable impedances. A fully connected RIS with $D$ reflecting elements operates as a $D$-port reciprocal network. The interaction among elements is captured by the scattering matrix $\bm{\Theta}\in\mathbb{C}^{D\times D}$, which characterizes the reflection behavior of the system. Energy conservation imposes the constraint $\bm{\Theta}\bm{\Theta}^{H}\preceq\mathbf{I}_{D}$, implying that the Frobenius norm satisfies $\frac{1}{\sqrt{D}}|\bm{\Theta}|_{F}\leq 1$. When the circuit is lossless, the scattering matrix becomes unitary, i.e., $\bm{\Theta}\bm{\Theta}^{H}=\mathbf{I}_{D}$. Furthermore, due to the reciprocal nature of the BD-RIS hardware, the symmetry condition $\bm{\Theta}=\bm{\Theta}^{T}$ must also be satisfied. This property implies that $\frac{D(D+1)}{2}$ independent coefficients are needed to fully define the entries of $\bm{\Theta}$.
The group-connected architecture introduces a balance between complexity and performance. In this setup, the reflecting elements are divided into $G$ independent groups, each of which forms a fully connected subnetwork. As a result, the overall scattering matrix exhibits a block diagonal structure:
$\bm{\Theta}=\mathrm{bdiag}(\bm{\Theta}_{1},\cdots,\bm{\Theta}_{G})\in\mathbb{C}^{D\times D},$ (1)
where each submatrix $\bm{\Theta}_{g}\in\mathbb{C}^{D_{g}\times D_{g}}$ models the signal reflections induced by its corresponding fully connected group, with $D_{g}$ denoting the number of elements of the $g$th group, such that $D=\sum_{g}^{G}D_{g}$. Similar to the fully connected case, each submatrix satisfies the conditions $\bm{\Theta}_{g}=\bm{\Theta}_{g}^{T}$ and $\bm{\Theta}_{g}\bm{\Theta}_{g}^{H}\preceq\mathbf{I}_{D_{g}}$. The number of independent parameters needed to configure each group is, therefore, $\frac{D_{g}(D_{g}+1)}{2}$, in which $D_{g}\leq D$. From this point onward, we consider lossless reflection models, i.e., we assume $\bm{\Theta}\bm{\Theta}^{H}=\mathbf{I}_{D}$ and $\bm{\Theta}_{g}\bm{\Theta}_{g}^{H}=\mathbf{I}_{D_{g}}$ for the fully and group-connected RIS, respectively. The impact of different adversarial strategies is analyzed in the following sections.
III-B RSMA Protocol
In RSMA, the BS partitions each user’s message into common and private parts. Then, all common parts are encoded into a single symbol $s^{c}$, while each private part, intended for a specific user $u$, is modulated into a private symbol $s^{p}_{u}$.
Next, based on the acquired CSI, the BS designs two linear precoders: $\mathbf{w}^{c}\in\mathbb{C}^{M}$ to deliver the common symbol $s^{c}$ to all users and $\mathbf{w}^{p}_{u}\in\mathbb{C}^{M}$ to transmit the private symbol $s^{p}_{u}$ to its intended user. Finally, the precoded symbols are superimposed in the power domain, resulting in the transmitted signal:
$\mathbf{s}$ $=\mathbf{w}^{c}\sqrt{P\alpha^{c}}s^{c}+\sum_{i\in\mathcal{U}}\mathbf{w}^{p}_{i}\sqrt{P\alpha_{i}^{p}}s_{u}^{p}\in\mathbb{C}^{M},$ (2)
where $\alpha^{c}$ and $\alpha_{i}^{p}$ are the power coefficients for the common and private symbols, respectively, satisfying $\alpha^{c}+\sum_{i\in\mathcal{U}}\alpha_{i}^{p}=1$, with $P$ denoting the total transmission power. Users receive the superimposed data streams through both the direct link and the adversarial BD-RIS reflections. Thus, the received signal at user $u$ is given by:
$r_{u}=\mathbf{g}_{u}^{H}\bm{\Theta}\mathbf{G}\mathbf{s}+\mathbf{h}_{u}^{H}\mathbf{s}+n_{u},$ (3)
where $\mathbf{h}_{u}\in\mathbb{C}^{M}$, $\mathbf{G}\in\mathbb{C}^{D\times M}$, and $\mathbf{g}_{u}\in\mathbb{C}^{D}$ model the propagation channels for the direct BS-user link, the BS-RIS link, and the RIS-user link, respectively, and $n_{u}\in\mathbb{C}$ denotes the additive white Gaussian noise with variance $\sigma^{2}$.
Each user first decodes the common message, while treating all private signals as interference. Consequently, the SINR (SINR) for the common message at user $u$ is:
$\gamma^{c}_{u}=\frac{|\big(\mathbf{g}_{u}^{H}\bm{\Theta}\mathbf{G}+\mathbf{h}_{u}^{H}\big)\mathbf{w}^{c}|^{2}P\alpha^{c}}{\sum_{i\in\mathcal{U}}|\big(\mathbf{g}_{u}^{H}\bm{\Theta}\mathbf{G}+\mathbf{h}_{u}^{H}\big)\mathbf{w}^{p}_{i}|^{2}P\alpha_{i}^{p}+\sigma^{2}}.$ (4)
Since all users must successfully decode the common message, the corresponding achievable rate is dictated by the weakest user, resulting in:
$R^{c}=\min_{u\in\mathcal{U}}R_{u}^{c},\quad\text{where}\quad R_{u}^{c}=\log_{2}(1+\gamma_{u}^{c}).$ (5)
Once the common message is recovered, SIC is applied to remove its contribution from $r_{u}$. However, due to factors such as hardware impairments and estimation errors, the user’s reconstructed common signal might not perfectly cancel the received one. This mismatch leads to imperfect SIC, which can be modeled as a linear function of the common signal power[28]:
$\chi^{c}_{u}=\xi\left|(\mathbf{g}_{u}^{H}\bm{\Theta}\mathbf{G}+\mathbf{h}_{u}^{H})\mathbf{w}^{c}\right|^{2}P\alpha^{c},$ (6)
where $\xi\in[0,1]$ denotes the SIC error factor. A smaller $\xi$ indicates a higher quality of SIC, with $\xi=0$ representing the ideal case of perfect cancellation. This residual interference is treated as additional noise when decoding the private message. Consequently, the private SINR at user $u$ is given by:
$\gamma^{p}_{u}=\frac{|\big(\mathbf{g}_{u}^{H}\bm{\Theta}\mathbf{G}+\mathbf{h}_{u}^{H}\big)\mathbf{w}^{p}_{u}|^{2}P\alpha_{u}^{p}}{\sum_{i\in\mathcal{U},i\neq u}|\big(\mathbf{g}_{u}^{H}\bm{\Theta}\mathbf{G}+\mathbf{h}_{u}^{H}\big)\mathbf{w}^{p}_{i}|^{2}P\alpha_{i}^{p}+\chi^{c}_{u}+\sigma^{2}}.$ (7)
The achievable private rate at user $u$ is then:
$R_{u}^{p}=\log_{2}(1+\gamma_{u}^{p}),$ (8)
yielding a system sum rate of:
$R=R^{c}+\sum_{i\in\mathcal{U}}R_{i}^{p}.$ (9)
III-C Quantifying Robustness of RSMA
We propose two performance metrics to evaluate the robustness of RSMA against RIS-induced attacks, namely rate degradation and robustness index. The rate degradation quantifies the performance loss caused by an attack, computed as the difference between the baseline performance in a secure environment and the degraded performance during an attack. To this end, the common rate is assumed to be uniformly allocated among the users, i.e., $R^{c}/U$. The overall effective rate for user $u$ is then the sum of this portion of the common rate and the individual private rate $R^{p}_{u}$. The rate degradation experienced by the $u$th user is defined as
$\Delta R_{u}=\bar{R}^{p}_{u}-\tilde{R}^{p}_{u}+\frac{\bar{R}^{c}-\tilde{R}^{c}}{U},$ (10)
where $\tilde{R}^{p}_{u}$ and $\tilde{R}^{c}$ are the private and common rates observed when the system is under attack, and $\bar{R}^{p}_{u}$ and $\bar{R}^{c}$ are the reference rates in the absence of the attack.
The robustness index $\kappa$ complements the rate degradation metric by evaluating the proportion of baseline performance retained during an attack. This normalized metric quantifies the system’s ability to maintain performance under RIS-induced attacks relative to its baseline rates $\bar{R}^{p}_{u}$ and $\bar{R}^{c}$. The robustness index is defined as
$\kappa=\frac{1}{U}\sum_{u=1}^{U}\left(1-\frac{\Delta R_{u}}{\frac{\bar{R}^{c}}{U}+\bar{R}^{p}_{u}}\right).$ (11)
Note that a higher degradation $\Delta R_{u}$ leads to a lower robustness index $\kappa\in[0,1]$, such that $\kappa=0$ indicates complete collapse, while $\kappa=1$ corresponds to complete robustness.
III-D CSI Acquisition
Under the assumption of reciprocity, the downlink channels can be estimated at the BS based on uplink measurements from pilot signals transmitted by the users. As explained, during this phase, the attacker can configure the BD-RIS to either absorb or randomly reflect incoming signals. In absorption mode, the pilot signals impinging on the BD-RIS are prevented from reaching the BS. In the random reflective mode, on the other hand, the pilot signals will characterize the BD-RIS-related propagation paths. However, the attacker reconfigures the reflection coefficients in the data transmission phase, making the acquired CSI inaccurate. Thus, in both modes, only imperfect estimates of the channels $\mathbf{h}_{u}$ are accessible to the BS. Adding to the BD-RIS distortions, other rapid changes in the propagation environment and hardware limitations can also introduce errors in the estimation process. As a result, the direct link estimates will be a combination of the true channels and error components, modeled as:
$\hat{\mathbf{h}}_{u}=\sqrt{1-\epsilon^{\text{\tiny BS-U}}}\,(\mathbf{h}_{u}+\psi\mathbf{G}^{H}\bm{\Theta}\mathbf{g}_{u})+\sqrt{\epsilon^{\text{\tiny BS-U}}}\,\mathbf{e}^{\text{\tiny BS-U}}_{u},$ (12)
where $\hat{\mathbf{h}}_{u}\in\mathbb{C}^{M}$ is the estimated channel, $\mathbf{e}^{\text{\tiny BS-U}}_{u}\in\mathbb{C}^{M}$ is a standard complex Gaussian error vector, $\psi$ is a binary coefficient modeling the operation mode of the BD-RIS, with $\psi=0$ corresponding to absorption mode, and $\psi=1$ corresponding to reflective mode, and $\epsilon^{\text{\tiny BS-U}}\in[0,1]$ is the error factor for the BS-user link. This model indicates that when $\epsilon^{\text{\tiny BS-U}}=0$, the channel estimate is perfect (i.e., $\hat{\mathbf{h}}_{u}=\mathbf{h}_{u}+\psi\mathbf{G}^{H}\bm{\Theta}\mathbf{g}_{u}$), while increasing values of $\epsilon^{\text{\tiny BS-U}}$ imply a larger error component in the estimate. Similarly, when $\psi=1$, the uplink CSI is further perturbed by the adversarial BD-RIS link.
As for the attacker, we assume that it can only estimate the channels associated with its BD-RIS, also modeled as imperfect. Specifically, the estimates of the BS-RIS channel matrix and the RIS-user channel vector are given by
$\hat{\mathbf{G}}=\sqrt{1-\epsilon^{\text{\tiny BS-RIS}}}\,\mathbf{G}+\sqrt{\epsilon^{\text{\tiny BS-RIS}}}\,\mathbf{E}^{\text{\tiny BS-RIS}},$ (13)
and
$\hat{\mathbf{g}}_{u}=\sqrt{1-\epsilon^{\text{\tiny RIS-U}}}\,\mathbf{g}_{u}+\sqrt{\epsilon^{\text{\tiny RIS-U}}}\,\mathbf{e}^{\text{\tiny RIS-U}}_{u},$ (14)
where $\mathbf{E}^{\text{\tiny BS-RIS}}$ and $\mathbf{e}^{\text{\tiny RIS-U}}_{u}$ are the corresponding complex Gaussian error components, with $\epsilon^{\text{\tiny BS-RIS}}$ and $\epsilon^{\text{\tiny RIS-U}}$ denoting the error factors for the BS-RIS and RIS-user links, respectively.
III-E BS Precoder Design
The private precoder should ensure that the interference observed at user $u$ due to transmissions intended for any other user $u^{\prime}$ is negligible, i.e., $\mathbf{h}_{u^{\prime}}^{H}\mathbf{w}_{u}^{p}\approx 0$, $\forall\,u^{\prime}\neq u$. This objective is achieved via a regularized zero-forcing approach, which effectively mitigates inter-user interference while balancing noise enhancement. Specifically, the BS first constructs the estimated direct-link channel matrix as
$\hat{\mathbf{H}}=\begin{bmatrix}\hat{\mathbf{h}}_{1},\hat{\mathbf{h}}_{2},\cdots,\hat{\mathbf{h}}_{U}\end{bmatrix}\in\mathbb{C}^{M\times U}.$ (15)
Then, assuming that $M\geq U$, which guarantees that $\hat{\mathbf{H}}$ has full column rank, the regularized zero-forcing precoding matrix can be computed as
$\mathbf{W}^{p}=\hat{\mathbf{H}}\left(\hat{\mathbf{H}}^{H}\hat{\mathbf{H}}+\omega\mathbf{I}\right)^{-1}\in\mathbb{C}^{M\times U},$ (16)
where the regularization factor is given by $\omega={\sigma^{2}}/{P}$. The private precoder for user $u$ can be readily obtained by
$\mathbf{w}_{u}^{p}=\beta^{p}_{u}{\left[\mathbf{W}^{p}\right]_{:,u}},$ (17)
where $\beta^{p}_{u}=\frac{1}{(\mathbf{w}_{u}^{p})^{H}\mathbf{w}_{u}^{p}}$ is the normalization factor for user $u$.
The common precoder $\mathbf{w}^{c}$ should reliably deliver the common message across all users. However, its optimal design is NP-hard, while approximate solutions typically require iterative methods with high computational complexity [29]. To tackle this challenge, we adopt a low-complexity weighted MF (MF) approach as follows:
$\mathbf{w}^{c}=\beta^{c}\sum_{i\in\mathcal{U}}\upsilon_{i}\hat{\mathbf{h}}_{i},$ (18)
where the normalization factor $\beta^{c}=\left\|\sum_{i\in\mathcal{U}}\upsilon_{i}\hat{\mathbf{h}}_{i}\right\|_{2}^{-1}$ guarantees that $\|\mathbf{w}^{c}\|_{2}=1$, and $\upsilon_{i}$ is the weight for the $i$th user. In particular, we prioritize users with weaker channel conditions by setting $\upsilon_{i}=1/{\|\hat{\mathbf{h}}_{i}\|_{2}^{2}}$. This strategy ensures that users experiencing weaker channels receive higher priority in the common precoder design.
III-F Power Allocation
The BS employs a power allocation policy aiming at the maximization of the system sum rate. Given the precoders $\mathbf{w}^{c}$ and $\mathbf{w}^{p}_{u}$ for each user $u\in\mathcal{U}$, the BS aims to solve the following optimization problem:
$\max_{\alpha^{c},\,\{\alpha_{u}^{p}\}_{u\in\mathcal{U}}}\quad$ $R^{c}+\sum_{u\in\mathcal{U}}R_{u}^{p},$ (19a)
s.t. $\alpha^{c}+\sum_{u\in\mathcal{U}}\alpha_{u}^{p}=1,$ (19b)
$\alpha_{u}^{p}\geq 0,\quad\alpha^{c}\geq 0.$ (19c)
A closed-form optimal solution is not possible due to the minimum operator in the common rate in (5), i.e., $R^{c}=\min_{u\in\mathcal{U}}R_{u}^{c}$, and the coupled power coefficients, which make the problem non-convex. Instead, we propose a simple but effective approach to approximate our objective.
A meaningful way to perform power allocation is to keep the interference resulting from imperfect CSI approximately at the same level as the noise floor [11]. By doing this, the degradation of users’ data rates can be mitigated. In our system, which considers imperfect SIC, the noise floor for the private messages is a combination of the thermal noise $\sigma^{2}$ and the residual SIC interference $\chi^{c}_{u}$. Therefore, we aim to ensure that the SINR expression for the private stream, $\gamma_{u}^{p}$, is on the same order as the sum of this effective noise power, i.e., $\sum_{i\neq u}|\mathbf{h}_{u}^{H}\mathbf{w}^{p}_{i}|^{2}P\alpha^{p}_{u}\approx\xi\left|\mathbf{h}_{u}^{H}\mathbf{w}^{c}\right|^{2}P\alpha^{c}+\sigma^{2}$. To achieve this goal, we first adopt a uniform power allocation across the private messages, i.e., $\alpha^{p}=\alpha^{p}_{u},\forall u\in\mathcal{U}$. Then, we can substitute $\alpha^{c}=1-U\alpha^{p}$ and show for the $u$th user that:
$\sum_{i\neq u}|\mathbf{h}_{u}^{H}\mathbf{w}^{p}_{i}|^{2}P\alpha^{p}\approx\xi\left|\mathbf{h}_{u}^{H}\mathbf{w}^{c}\right|^{2}P\alpha^{c}+\sigma^{2}$
$\implies\alpha^{p}\propto\frac{\xi|\mathbf{h}_{u}^{H}\mathbf{w}^{c}|^{2}P+\sigma^{2}}{\sum_{i\neq u}|\mathbf{h}_{u}^{H}\mathbf{w}^{p}_{i}|^{2}P+\xi U|\mathbf{h}_{u}^{H}\mathbf{w}^{c}|^{2}P}.$ (20)
That is, the desired $\alpha^{p}$ should be directly proportional to the effective noise floor and inversely proportional to the private inter-user interference.
Note that we need to compute an $\alpha^{p}$ capable of balancing the interference terms across all users in such a way that the system’s sum rate is maximized. Given this goal, we prioritize the user with the best channel conditions, i.e., the one that contributes the most to the sum rate. To this end, we select as reference the best performance indicator given by:
$u^{\star}=\arg\max_{u\in\mathcal{U}}\hskip 5.69054pt\frac{|\mathbf{h}_{u}^{H}\mathbf{w}^{p}_{u}|^{2}P}{\sum_{i\neq u}|\mathbf{h}_{u}^{H}\mathbf{w}^{p}_{i}|^{2}P+\xi|\mathbf{h}_{u}^{H}\mathbf{w}^{c}|^{2}P}+\sigma^{2}.$ (21)
After determining the reference user, we recall (20) and transform the allocation into a tractable problem by introducing a scaling factor $\eta$, as follows:
$\alpha^{p}=\frac{\eta\left(\xi|\mathbf{h}_{u^{\star}}^{H}\mathbf{w}^{c}|^{2}P+\sigma^{2}\right)}{\sum_{i\neq u^{\star}}|\mathbf{h}_{u^{\star}}^{H}\mathbf{w}^{p}_{i}|^{2}P+\xi U|\mathbf{h}_{u^{\star}}^{H}\mathbf{w}^{c}|^{2}P}.$ (22)
The common power coefficient is consequently given by $\alpha^{c}=1-U\alpha^{p}$. The scaling factor $\eta$ adjusts the aggressiveness of the allocation, i.e., a higher $\eta$ increases private power allocation, while a lower $\eta$ reserves more power for the common message.
Thus, the original formulation can be transformed into a simpler version with a single optimization variable, $\eta$, resulting in the following relaxed power allocation problem:
$\max_{\eta}\quad$ $R^{c}(\eta)+\sum_{u\in\mathcal{U}}R_{u}^{p}(\eta),$ (23a)
s.t. $\alpha^{p}=\frac{\eta\left(\xi|\mathbf{h}_{u^{\star}}^{H}\mathbf{w}^{c}|^{2}P+\sigma^{2}\right)}{\sum_{i\neq u^{\star}}|\mathbf{h}_{u^{\star}}^{H}\mathbf{w}^{p}_{i}|^{2}P+\xi U|\mathbf{h}_{u^{\star}}^{H}\mathbf{w}^{c}|^{2}P},$ (23b)
$\alpha^{c}=1-U\alpha^{p},$ (23c)
$U\alpha^{p}\geq 0,\quad\alpha^{c}\geq 0.$ (23d)
where $R^{c}(\eta)$ and $R_{u}^{p}(\eta)$ denote the common and private rates as functions of $\eta$. Although the problem remains non-convex due to the minimum operator in the common rate and the nonlinear dependence of the SINRs on $\eta$, a linear search over $\eta$ within its feasible range provides a practical and efficient method to approximate the solution of (23). Specifically, if $\eta$ is chosen such that $0\leq\eta\leq\frac{\sum_{i\neq u^{\star}}|\mathbf{h}_{u^{\star}}^{H}\mathbf{w}^{p}_{i}|^{2}P+\xi U|\mathbf{h}_{u^{\star}}^{H}\mathbf{w}^{c}|^{2}P}{U\left(\xi|\mathbf{h}_{u^{\star}}^{H}\mathbf{w}^{c}|^{2}P+\sigma^{2}\right)}$, the constraints in (23d) are automatically satisfied.
IV Design of BD-RIS-Induced Attacks
The malicious goal of the attacker is to compute a reflection matrix $\bm{\Theta}$ that degrades the performance of the employed RSMA scheme. In the following subsections, we investigate two approaches to accomplish the goal.
IV-A Random Interference Attack
A simple yet effective method to degrade the performance of RSMA involves randomly configuring the reflection coefficients of an adversarial RIS to introduce interference, as demonstrated in our previous work with diagonal RIS [6].
Fully Connected Architecture
As explained, a fully connected RIS requires its reflection matrix to be both symmetric and unitary. To obtain a valid set of reflection coefficients, the adversary starts by creating a random complex matrix $\tilde{\bm{\Gamma}}\in\mathbb{C}^{D\times D}$ and then symmetrizes it, as follows:
$\bm{\Gamma}$ $=\frac{1}{2}\left(\tilde{\bm{\Gamma}}+\tilde{\bm{\Gamma}}^{T}\right).$ (24)
However, $\bm{\Gamma}$ is generally not unitary. Therefore, we need to project it onto the set of symmetric unitary matrices while respecting its symmetry constraint. To this end, we can rely on the Takagi factorization, which is introduced next.
Property I
Let $\bm{\Gamma}$ be a symmetric complex matrix. Then, the Takagi factorization states that there exists a unitary matrix $\mathbf{U}_{\bm{\Gamma}}$ and a diagonal matrix $\bm{\Sigma}_{\bm{\Gamma}}$ such that $\bm{\Gamma}=\mathbf{U}_{\bm{\Gamma}}\bm{\Sigma}_{\bm{\Gamma}}\mathbf{U}_{\bm{\Gamma}}^{T}$, where the diagonal entries of $\bm{\Sigma}_{\bm{\Gamma}}$ are nonnegative singular values of $\bm{\Gamma}$, called Takagi values, and the columns of $\mathbf{U}_{\bm{\Gamma}}$ are the associated orthonormal Takagi vectors. With Property IV-A, the desired symmetric unitary reflection matrix can be achieved through the following lemma.
Lemma I
Let $\bm{\Gamma}$ be a complex symmetric matrix that admits the Takagi factorization $\bm{\Gamma}=\mathbf{U}_{\bm{\Gamma}}\mathbf{\Sigma}_{\bm{\Gamma}}\mathbf{U}_{\bm{\Gamma}}^{T}$. Then, the unique symmetric unitary matrix
$\mathbf{\Theta}=\mathbf{U}_{\bm{\Gamma}}\mathbf{U}_{\bm{\Gamma}}^{T},$ (25)
minimizes the Frobenius norm $\|\bm{\Gamma}-\mathbf{\Theta}\|_{F}$ and, thus, defines the optimal projection onto the set of symmetric unitary matrices.
Proof: Since $\bm{\Gamma}$ is complex symmetric, its Takagi factorization is given by $\bm{\Gamma}=\mathbf{U}_{\bm{\Gamma}}\mathbf{\Sigma}_{\bm{\Gamma}}\mathbf{U}_{\bm{\Gamma}}^{T}$. Any symmetric unitary matrix can be written in the form $\mathbf{\Theta}=\mathbf{U}_{\bm{\Gamma}}\mathbf{D}\mathbf{U}_{\bm{\Gamma}}^{T}$, where $\mathbf{D}$ is a diagonal matrix with entries on the unit circle, i.e., $|[\mathbf{D}]_{ii}|=1$, for each diagonal element. The problem of finding the symmetric unitary matrix closest to $\bm{\Gamma}$ reduces to choosing $\mathbf{D}$ such that
$\|\bm{\Gamma}-\mathbf{\Theta}\|_{F}$ $=\|\mathbf{U}_{\bm{\Gamma}}\mathbf{\Sigma}_{\bm{\Gamma}}\mathbf{U}_{\bm{\Gamma}}^{T}-\mathbf{U}_{\bm{\Gamma}}\mathbf{D}\mathbf{U}_{\bm{\Gamma}}^{T}\|_{F}$
$=\|\mathbf{\Sigma}_{\bm{\Gamma}}-\mathbf{D}\|_{F}$ (26)
is minimized, where the last equality follows from the unitary invariance of the Frobenius norm. Since each diagonal entry $\bm{\Sigma}_{\bm{\Gamma}}]_{ii}$ is nonnegative, for each $i$ the distance $\left|[\bm{\Sigma}_{\bm{\Gamma}}]_{ii}-[\mathbf{D}]_{ii}\right|$ is minimized by choosing $\mathbf{D}]_{ii}=1$, as $\left|[\bm{\Sigma}_{\bm{\Gamma}}]_{ii}-e^{i\theta}\right|$ attains its minimum when $\theta=0$. Therefore, the optimal projection is given by $\mathbf{D}=\mathbf{I}$, yielding $\mathbf{\Theta}=\mathbf{U}_{\bm{\Gamma}}\mathbf{I}\mathbf{U}_{\bm{\Gamma}}^{T}=\mathbf{U}_{\bm{\Gamma}}\mathbf{U}_{\bm{\Gamma}}^{T}$, which, clearly, is symmetric since
$\mathbf{\Theta}^{T}=(\mathbf{U}_{\bm{\Gamma}}\mathbf{U}_{\bm{\Gamma}}^{T})^{T}=\mathbf{U}_{\bm{\Gamma}}\mathbf{U}_{\bm{\Gamma}}^{T}=\mathbf{\Theta}.$ (27)
This completes the proof. $\square$
The projection through the Takagi factorization respects both the symmetry and unitary constraints required by the fully connected RIS. For any invertible complex symmetric matrix with distinct singular values, the Takagi factorization can be implemented very efficiently by aligning the phases of the singular vectors computed via the SVD [30].
Group-Connected Architecture
A similar strategy can be extended for a group-connected RIS architecture. In this case, the attacker must respect the group-wise constraints. Specifically, the reflection matrix $\mathbf{\Theta}$ should be block-diagonal, comprising $G$ independent sub-matrices $\mathbf{\Theta}_{g}$, for $g=1,\cdots,G$, each one being symmetric and unitary. To generate such a matrix, the adversary follows the same process as for the fully connected RIS, but now for each group. First, it constructs random complex matrices $\tilde{\mathbf{\Gamma}}_{g}\in\mathbb{C}^{D_{g}\times D_{g}}$ for every group. Then, symmetrized versions $\mathbf{\Gamma}_{g}$ are achieved as in (24). Subsequently, the Takagi factorization is computed for each symmetrized matrix, as $\mathbf{\Gamma}_{g}=\mathbf{U}_{\mathbf{\Gamma}_{g}}\mathbf{\Sigma}_{\mathbf{\Gamma}_{g}}\mathbf{U}_{\mathbf{\Gamma}_{g}}^{T}$. The symmetric unitary matrix for the $g$th group is then obtained via the projection $\mathbf{\Theta}_{g}=\mathbf{U}_{\mathbf{\Gamma}_{g}}\mathbf{U}_{\mathbf{\Gamma}_{g}}^{T}$, as in Lemma IV-A. Finally, the full reflection matrix is assembled by organizing the $G$ sub-matrices into a block-diagonal structure, i.e., $\mathbf{\Theta}=\mathrm{bdiag}(\mathbf{\Theta}_{1},\ldots,\mathbf{\Theta}_{G})$.
IV-B Aligned interference attack
A more sophisticated adversarial strategy is to optimize the BD-RIS reflection coefficients to maximize the interference induced at the users. As demonstrated in [6] for conventional RIS systems, if the adversary can acquire partial RIS-associated CSI [^1] To acquire the necessary CSI, an adversary could exploit semi-passive RIS architectures, where active sensing elements are incorporated to estimate the RIS channels over time [31]., it becomes possible to strategically steer the reflected power toward the users. In this subsection, we generalize this strategy to BD-RIS architectures.
Fully Connected Architecture
Let $\mu_{i}\in(0,1]$ denote the adversarial weight assigned to user $i\in\mathcal{U}$. The malicious fully connected RIS controller aims to maximize the weighted sum of interference powers received at users, subject to unitary and symmetry constraints. Formally, the optimization problem can be formulated as:
$\max_{\bm{\Theta}}\quad$ $\sum_{i\in\mathcal{U}}\mu_{i}\left\|\mathbf{g}_{i}^{H}\bm{\Theta}\mathbf{G}\right\|_{2}^{2},$ (28a)
s.t. $\bm{\Theta}=\bm{\Theta}^{T}.$ (28b)
$\bm{\Theta}^{H}\bm{\Theta}=\mathbf{I},$ (28c)
The unitary constraint in (28c) defines a non-convex Stiefel manifold. This, combined with the quadratic dependence of the objective function on $\bm{\Theta}$, makes the problem challenging to solve optimally. To simplify the original problem, we start by addressing the symmetry constraint in (28b). To this end, we exploit the mathematical structure of symmetric matrices by vectorizing $\bm{\Theta}$ using the concept of duplication matrix.
Property II
For any symmetric matrix $\bm{\Theta}$, there exists a duplication matrix $\mathbf{D}\in\mathbb{R}^{D^{2}\times\frac{D(D+1)}{2}}$ that maps the half-vectorization of $\bm{\Theta}$ to its full vectorization:
$\mathrm{vec}(\bm{\Theta})=\mathbf{D}\mathrm{vech}(\bm{\Theta}),$ (29)
where the duplication matrix $\mathbf{D}$ is uniquely defined as a sparse matrix with binary entries, given by:
$\mathbf{D}^{T}\triangleq\sum_{i=1}^{D}\sum_{j=1}^{i}\bm{u}_{ij}\mathrm{vec}(\mathbf{T}_{ij})^{T},$ (30)
where $\bm{u}_{ij}$ is a unit vector of length $\frac{D(D+1)}{2}$ with a single non-zero entry at position $(j-1)D+i-\frac{j(j-1)}{2}]$, and $\mathbf{T}_{ij}$ is a $D\times D$ matrix with ones at positions $(i,j)$ and $(j,i)$, and zeros elsewhere [32, Definitions 3.2a and 3.2b].
Property IV-B inherently satisfies the symmetry constraint of the original formulation, allowing us to reformulate the problem in a simpler relaxed version in terms of $\mathrm{vech}(\bm{\Theta})$.
Input: Channel matrix $\mathbf{G}$, user channels $\{\mathbf{g}_{i}\}_{i\in\mathcal{U}}$, duplication matrix $\mathbf{D}$, user weights $\{\mu_{i}\}_{i\in\mathcal{U}}$
Output: Optimized reflection matrix $\bm{\Theta}^{*}$
1
2
3Compute the transformed interference channel matrices:
4foreach $i\in\mathcal{U}$ do
5   Compute: $\mathbf{S}_{i}\leftarrow(\mathbf{G}^{T}\otimes\mathbf{g}_{i}^{H})$;
6
7Construct the weighted interference matrix:
$\mathbf{S}\leftarrow[\sqrt{\mu_{1}}\mathbf{S}_{1}^{H},\cdots,\sqrt{\mu_{U}}\mathbf{S}_{U}^{H}]^{H}.$
8Compute the SVD of $\bar{\bm{S}}=\mathbf{S}\mathbf{D}$ to solve the relaxed problem in (33):
$\bar{\bm{S}}=\mathbf{U}_{\bar{\bm{S}}}\bm{\Sigma}_{\bar{\bm{S}}}\mathbf{V}_{\bar{\bm{S}}}^{H}.$
9Obtain the optimized reduced coefficient vector: $\bm{\theta}\leftarrow\mathbf{v}_{\bar{\bm{S}},1}$, where $\mathbf{v}_{\bar{\bm{S}},1}$ is the corresponding dominant right singular vector.
10Construct the relaxed reflection matrix: $\bm{\Phi}\leftarrow\mathrm{unvec}(\mathbf{D}\bm{\theta})$.
11Perform the Takagi factorization: $\bm{\Phi}=\mathbf{U}_{\bm{\Phi}}\bm{\Sigma}_{\bm{\Phi}}\mathbf{U}_{\bm{\Phi}}^{T}$.
12Project onto the set of unitary matrices: $\bm{\Theta}^{*}\leftarrow\mathbf{U}_{\bm{\Phi}}\mathbf{U}_{\bm{\Phi}}^{T}$.
return $\bm{\Theta}^{*}$.
Algorithm 1 Aligned Interference Attack for Fully Connected RIS
Using the Kronecker product identity $\mathrm{vec}(\mathbf{A}\bm{\Theta}\mathbf{C})=(\mathbf{C}^{T}\otimes\mathbf{A})\mathrm{vec}(\bm{\Theta})$, and invoking the duplication matrix relation in (29), we can vectorize the interference channels $\mathbf{g}_{i}^{H}\bm{\Theta}\mathbf{G}$ as:
$\mathrm{vec}(\mathbf{g}_{i}^{H}\bm{\Theta}\mathbf{G})$ $=(\mathbf{G}^{T}\otimes\mathbf{g}_{i}^{H})\mathbf{D}\mathrm{vech}(\bm{\Theta}).$ (31)
Then, by defining $\mathbf{S}_{i}=(\mathbf{G}^{T}\otimes\mathbf{g}_{i}^{H})\in\mathbb{C}^{M\times D^{2}}$, and letting $\bm{\theta}=\mathrm{vech}(\bm{\Theta})\in\mathbb{C}^{\frac{D(D+1)}{2}}$ denote the reduced vector of unique reflection coefficients, the objective function in (28a) becomes:
$\sum_{i\in\mathcal{U}}\mu_{i}$ $\left\|(\mathbf{G}^{T}\otimes\mathbf{g}_{i}^{H})\mathbf{D}\bm{\theta}\right\|_{2}^{2}$
$=\sum_{i\in\mathcal{U}}\mu_{i}\left\|\mathbf{S}_{i}\mathbf{D}\bm{\theta}\right\|_{2}^{2}$
$=\bm{\theta}^{H}\mathbf{D}^{T}\left(\sum_{i\in\mathcal{U}}\mu_{i}\mathbf{S}_{i}^{H}\mathbf{S}_{i}\right)\mathbf{D}\bm{\theta}.$ (32)
Input: Channel matrix $\mathbf{G}$, user channels $\{\mathbf{g}_{i}\}_{i\in\mathcal{U}}$, RIS group indices $\{\mathcal{E}_{g}\}_{g\in\mathcal{G}}$, duplication matrices $\{\mathbf{D}_{g}\}_{g\in\mathcal{G}}$, user weights $\{\mu_{i}\}_{i\in\mathcal{U}}$
Output: Optimized block-diagonal reflection matrix $\bm{\Theta}^{*}$
1
2Preprocess each RIS group:
3foreach $g\in\mathcal{G}$ do
4 $D_{g}\leftarrow|\mathcal{E}_{g}|$;
5   Extract BS-RIS subchannel: $\mathbf{G}_{g}\leftarrow[\mathbf{G}]_{\mathcal{E}_{g},:}$;
6 foreach $i\in\mathcal{U}$ do
7      Extract RIS-user subchannel: $\mathbf{g}_{ig}\leftarrow[\mathbf{g}_{i}]_{\mathcal{E}_{g}}$;
8      Compute: $\mathbf{S}_{ig}\leftarrow\bigl(\mathbf{G}_{g}^{T}\otimes\mathbf{g}_{ig}^{H}\bigr)$;
9
10
11
12Construct the interference matrix from group contributions:
13foreach $i\in\mathcal{U}$ do
14   Compute the block-row matrix:
$\mathbf{J}_{i}\leftarrow\begin{bmatrix}\mathbf{S}_{i1}\mathbf{D}_{1}&\mathbf{S}_{i2}\mathbf{D}_{2}&\cdots&\mathbf{S}_{iG}\mathbf{D}_{G}\end{bmatrix}\,$.
15
16Construct the stacked global matrix:
$\mathbf{J}\leftarrow\begin{bmatrix}\sqrt{\mu_{1}}\mathbf{J}_{1}^{T},\sqrt{\mu_{2}}\mathbf{J}_{2}^{T},\cdots,\sqrt{\mu_{U}}\mathbf{J}_{U}^{T}\end{bmatrix}^{T}\,.$
17Compute the SVD of $\mathbf{J}$ to solve the relaxed problem in (44):
$\mathbf{J}=\mathbf{U}_{\mathbf{J}}\bm{\Sigma}_{\mathbf{J}}\mathbf{V}_{\mathbf{J}}^{H}\,.$
18Obtain the optimized coefficient vector: $\bm{\varphi}\leftarrow\mathbf{v}_{\mathbf{J},1}$, where $\mathbf{v}_{\mathbf{J},1}$ is the corresponding dominant right singular vector.
19foreach $g\in\mathcal{G}$ do
20   Extract reduced vector: $\bm{\theta}_{g}\leftarrow[\bm{\varphi}]_{\mathcal{R}_{g}}$, where $\mathcal{R}_{g}\subset\left\{1,\cdots,\sum_{g\in\mathcal{G}}\frac{D_{g}(D_{g}+1)}{2}\right\}$ is the index set for group $g$, with $|\mathcal{R}_{g}|=\frac{D_{g}(D_{g}+1)}{2}$.
21  Construct the relaxed reflection matrix: $\bm{\Phi}_{g}\leftarrow\mathrm{unvec}(\mathbf{D}_{g}\,\bm{\theta}_{g})$.
22  Perform the Takagi factorization: $\bm{\Phi}_{g}=\mathbf{U}_{\bm{\Phi}_{g}}\bm{\Sigma}_{\bm{\Phi}_{g}}\mathbf{U}_{\bm{\Phi}_{g}}^{T}$.
23  Project onto the set of unitary matrices: $\bm{\Theta}^{*}_{g}\leftarrow\mathbf{U}_{\bm{\Phi}_{g}}\mathbf{U}_{\bm{\Phi}_{g}}^{T}.$
24
25Construct the final block-diagonal reflection matrix:
$\bm{\Theta}^{*}\leftarrow\mathrm{bdiag}\bigl(\bm{\Theta}^{*}_{1},\,\bm{\Theta}^{*}_{2},\,\dots,\,\bm{\Theta}^{*}_{G}\bigr).$
return $\bm{\Theta}^{*}$.
Algorithm 2 Aligned Interference Attack for Group-Connected RIS
Given that $\sum_{i\in\mathcal{U}}\mathbf{S}_{i}^{H}\mathbf{S}_{i}=\left[\mathbf{S}_{1}^{H},\cdots,\mathbf{S}_{U}^{H}\right]^{H}\left[\mathbf{S}_{1}^{H},\cdots,\mathbf{S}_{U}^{H}\right]$, we can define the tall weighted interference matrix: $\mathbf{S}\triangleq[\sqrt{\mu_{1}}\mathbf{S}_{1}^{H},\cdots,\sqrt{\mu_{U}}\mathbf{S}_{U}^{H}]^{H}\in\mathbb{C}^{UN\times D^{2}}$ and rewrite the expression in (IV-B) as the convex quadratic objective $\bm{\theta}^{H}\mathbf{D}^{T}\mathbf{S}^{H}\mathbf{S}\mathbf{D}\bm{\theta}$. This vectorized reformulation automatically satisfies the symmetry constraint in (28b) by construction. On the other hand, the non-convex unitary constraint in (28c) is relaxed with a convex $\ell_{2}$-norm bound $\|\bm{\theta}\|_{2}\leq 1$, which allow us to formulate a tractable optimization problem:
$\max_{\bm{\theta}}\quad$ $\bm{\theta}^{H}\mathbf{D}^{T}\mathbf{S}^{H}\mathbf{S}\mathbf{D}\bm{\theta}$ (33a)
s.t. $\|\bm{\theta}\|_{2}\leq 1.$ (33b)
This relaxed problem (33) constitutes a QCQP, whose optimal solution is the dominant right singular vector of the matrix $\bar{\bm{S}}\triangleq\mathbf{S}\mathbf{D}$. More specifically, by recalling the SVD, we can express $\bar{\bm{S}}=\mathbf{U}_{\bar{\bm{S}}}\bm{\Sigma}_{\bar{\bm{S}}}\mathbf{V}_{\bar{\bm{S}}}^{H}$, where $\mathbf{V}_{\bar{\bm{S}}}=[\mathbf{v}_{\bar{\bm{S}},1},\mathbf{v}_{\bar{\bm{S}},2},\dots,\mathbf{v}_{\bar{\bm{S}},\frac{D(D+1)}{2}}]$ contains the right singular vectors. Then, the optimal solution to the relaxed problem is given by $\bm{\theta}=\mathbf{v}_{\bar{\bm{S}},1}$, thus satisfying $\|\bm{\theta}\|_{2}=1$, and the corresponding relaxed BD-RIS matrix is obtained by:
$\bm{\Phi}=\mathrm{unvec}(\mathbf{D}\bm{\theta}).$ (34)
However, $\bm{\Phi}$ generally violates the unitary constraint, i.e., $\bm{\Phi}^{H}\bm{\Phi}\neq\mathbf{I}$. To enforce this, we project $\bm{\Phi}$ onto the closest unitary matrix via the Takagi factorization explained in Property IV-A, i.e., $\bm{\Phi}=\mathbf{U}_{\bm{\Phi}}\bm{\Sigma}_{\bm{\Phi}}\mathbf{U}_{\bm{\Phi}}^{T}$, allowing us to compute
$\bm{\Theta}^{*}=\mathbf{U}_{\bm{\Phi}}\mathbf{U}_{\bm{\Phi}}^{T},$ (35)
where, according to Lemma IV-A, the symmetry is preserved.
Group-Connected Architecture
The aligned interference attack is now generalized for the group-connected architecture. Let us consider a BD-RIS partitioned into $G$ independent groups, i.e., $\bm{\Theta}=\mathrm{bdiag}(\bm{\Theta}_{1},\dots,\bm{\Theta}_{G})$, with each block satisfying $\bm{\Theta}_{g}=\bm{\Theta}_{g}^{T}$ and $\bm{\Theta}_{g}^{H}\bm{\Theta}_{g}=\mathbf{I}$. Then, the optimized adversary attack strategy can be formulated as:
$\max_{\bm{\Theta}}\quad$ $\sum_{i\in\mathcal{U}}\mu_{i}\left\|\mathbf{g}_{i}^{H}\bm{\Theta}\mathbf{G}\right\|_{2}^{2},$ (36a)
s.t. $\bm{\Theta}=\mathrm{bdiag}(\bm{\Theta}_{1},\dots,\bm{\Theta}_{G}),$ (36b)
$\bm{\Theta}_{g}=\bm{\Theta}_{g}^{T},$ (36c)
$\bm{\Theta}_{g}^{H}\bm{\Theta}_{g}=\mathbf{I}.$ (36d)
To address problem 36, the first step is to decompose the global channel matrices into group-specific subchannels. To this end, let $\mathcal{E}_{g}\subset\{1,\dots,D\}$ denote the index subset of RIS elements associated with group $g$, with $|\mathcal{E}_{g}|=D_{g}$. Then, the BS-RIS subchannel for group $g$ can be expressed as
$\mathbf{G}_{g}=[\mathbf{G}]_{\mathcal{E}_{g},:}\in\mathbb{C}^{D_{g}\times M},$ (37)
while the RIS-user subchannel becomes
$\mathbf{g}_{ig}=[\mathbf{g}_{i}]_{\mathcal{E}_{g}}\in\mathbb{C}^{D_{g}}.$ (38)
Next, following Property IV-B, we construct a duplication matrix $\mathbf{D}_{g}\in\mathbb{R}^{D_{g}^{2}\times\frac{D_{g}(D_{g}+1)}{2}}$ to vectorize the symmetric matrix as $\mathrm{vec}(\bm{\Theta}_{g})=\mathbf{D}_{g}\bm{\theta}_{g}$ for each group, where $\bm{\theta}_{g}=\mathrm{vech}(\bm{\Theta}_{g})\in\mathbb{C}^{\frac{D_{g}(D_{g}+1)}{2}}$ denotes the $g$th reduced coefficient vector. We also define $\mathbf{S}_{ig}=(\mathbf{G}_{g}^{T}\otimes\mathbf{g}_{ig}^{H})\in\mathbb{C}^{M\times D_{g}^{2}}$, and represent the set of RIS groups by $\mathcal{G}=\{1,\cdots,G\}$. Then, we rewrite the adversary objective as follows:
$\sum_{i\in\mathcal{U}}\mu_{i}$ $\left\|\sum_{j\in\mathcal{G}}(\mathbf{G}_{j}^{T}\otimes\mathbf{g}_{ij}^{H})\mathbf{D}_{j}\bm{\theta}_{j}\right\|_{2}^{2}$
$\hskip 28.45274pt=\sum_{i\in\mathcal{U}}\mu_{i}\left\|\sum_{j\in\mathcal{G}}\mathbf{S}_{ij}\mathbf{D}_{j}\bm{\theta}_{j}\right\|_{2}^{2}.$ (39)
To further simplify, we introduce a global stacked vector:
$\bm{\varphi}\triangleq[\bm{\theta}_{1}^{T},\bm{\theta}_{2}^{T},\cdots,\bm{\theta}_{G}^{T}]^{T}\in\mathbb{C}^{\sum_{g\in\mathcal{G}}\frac{D_{g}(D_{g}+1)}{2}},$ (40)
and define, for each user $i\in\mathcal{U}$, the block row matrix:
$\mathbf{J}_{i}\triangleq\begin{bmatrix}\mathbf{S}_{i1}\mathbf{D}_{1},\mathbf{S}_{i2}\mathbf{D}_{2},\cdots,\mathbf{S}_{iG}\mathbf{D}_{G}\end{bmatrix}$. (41)
Now, the summation related to the groups can be written as a single matrix structure, as follows:
$\sum_{i\in\mathcal{U}}\mu_{i}$ $\left\|\sum_{j\in\mathcal{G}}\mathbf{S}_{ij}\mathbf{D}_{j}\bm{\theta}_{j}\right\|_{2}^{2}$
$=\bm{\varphi}^{H}\left(\sum_{i\in\mathcal{U}}\mu_{i}\mathbf{J}_{i}^{H}\mathbf{J}_{i}\right)\bm{\varphi}.$ (42)
Then, we define a global matrix $\mathbf{J}$ stacking all $\mathbf{J}_{i}$ matrices:
$\mathbf{J}\triangleq\begin{bmatrix}\sqrt{\mu_{1}}\mathbf{J}_{1}^{T},\sqrt{\mu_{2}}\mathbf{J}_{2}^{T},\cdots,\sqrt{\mu_{U}}\mathbf{J}_{U}^{T}\end{bmatrix}^{T}.$ (43)
This allows us to write the summation in (IV-B) in its matrix-equivalent form, resulting in the final compact problem:
$\underset{{\bm{\varphi}}}{\max}\hskip 8.53581pt$ $\bm{\varphi}^{H}\mathbf{J}^{H}\mathbf{J}\bm{\varphi},$ (44a)
s.t. $\|{\bm{\varphi}}\|_{2}\leq 1.$ (44b)
The solution to (44) can be obtained with the aid of the SVD, similarly to the problem (33). More specifically, we can decompose $\mathbf{J}=\mathbf{U}_{\mathbf{J}}\bm{\Sigma}_{\mathbf{J}}\mathbf{V}_{\mathbf{J}}^{H}$ and achieve the desired solution by computing $\bm{\varphi}=\mathbf{v}_{\mathbf{J},1}$, where $\mathbf{v}_{\mathbf{J},1}$ is the dominant right singular vector of $\mathbf{J}$, i.e., the first column of $\mathbf{V}_{\mathbf{J}}$. Finally, by letting $\mathcal{R}_{g}\subset\left\{1,\cdots,\sum_{g\in\mathcal{G}}\frac{D_{g}(D_{g}+1)}{2}\right\}$ denote the index subset corresponding to the $g$th reduced coefficient vector, with $|\mathcal{R}_{g}|=\frac{D_{g}(D_{g}+1)}{2}$, the relaxed coefficients for each independent RIS group can be obtained as $\bm{\theta}_{g}=[\bm{\varphi}]_{\mathcal{R}_{g}}$, and $\bm{\Phi}_{g}=\mathrm{unvec}(\mathbf{D}_{g}\bm{\theta}_{g})$, $\forall g\in\mathcal{G}$. Recalling Property IV-A, we compute the Takagi factorization $\bm{\Phi}_{g}=\mathbf{U}_{\bm{\Phi}_{g}}\bm{\Sigma}_{\bm{\Phi}_{g}}\mathbf{U}_{\bm{\Phi}_{g}}^{T}$ and project each $\bm{\Phi}_{g}$ onto its closest unitary matrix, as follows:
$\bm{\Theta}^{*}_{g}=\mathbf{U}_{\bm{\Phi}_{g}}\mathbf{U}_{\bm{\Phi}_{g}}^{T}.$ (45)
Finally, the full BD-RIS reflection matrix is constructed as $\bm{\Theta}^{*}=\mathrm{bdiag}(\bm{\Theta}^{*}_{1},\cdots,\bm{\Theta}^{*}_{G})$.
IV-C Complexity Analysis
In this subsection, we analyze the computational complexity of the proposed BD-RIS-induced attacks, in terms of complex FLOP, quantifying their scaling with key system parameters, such as the number of RIS elements $D$, transmit antennas $M$, and users $U$.
Random Interference Attack
The complexity load of this strategy is dominated by the generation of valid reflection matrices, as explained in Section IV-A. For the fully connected RIS, the procedure begins by generating and symmetrizing a $D\times D$ complex matrix $\mathbf{\Gamma}$, which requires $\mathcal{O}(D^{2})$ complex FLOP. The dominant cost is the projection step, performed via the Takagi factorization. This is implemented through the SVD of $\mathbf{\Gamma}$. Specifically, the standard SVD of an $A\times B$ matrix has complexity $\mathcal{O}(AB\min(A,B))$ [33]. Since $\mathbf{\Gamma}$ is $D\times D$, the associated cost is $\mathcal{O}(D^{3})$. The subsequent matrix multiplication, $\mathbf{\Theta}=\mathbf{U}_{\bm{\Gamma}}\mathbf{U}_{\bm{\Gamma}}^{T}$, also requires $\mathcal{O}(D^{3})$. Therefore, the total complexity is $\mathcal{O}(D^{2}+D^{3}+D^{3})$, yielding a worst-case complexity of $\mathcal{O}(D^{3})$.
For the group-connected RIS, the same procedure is carried out independently for each group. For the $g$th group, the generation and symmetrization of $\mathbf{\Gamma}_{g}$ requires $\mathcal{O}(D_{g}^{2})$, while the subsequent SVD and the final matrix multiplication require $\mathcal{O}(D_{g}^{3})$. The total complexity is therefore $\sum_{g=1}^{G}\mathcal{O}(D_{g}^{2}+D_{g}^{3}+D_{g}^{3})$. In the case where groups are equal-sized, i.e., $D_{g}=D/G$, the worst-case complexity becomes $\mathcal{O}(D^{3}/G^{2})$.
Aligned Interference Attack
Algorithm 1, for the fully connected RIS, begins by constructing the interference matrices (lines 1–4), which involves the computation of $U$ Kronecker products, yielding $\mathcal{O}(UMD^{2})$ complex FLOP. The subsequent multiplication $\bar{\mathbf{S}}=\mathbf{S}\mathbf{D}$ (line 5) can also be executed in $\mathcal{O}(UMD^{2})$ FLOP when exploiting the sparse structure of $\mathbf{D}$. The most expensive operation is the SVD of $\bar{\mathbf{S}}\in\mathbb{C}^{UM\times{D(D+1)}/{2}}$. When the RIS is relatively small, such that $D(D+1)/2<UM$, the matrix $\bar{\mathbf{S}}$ is tall and its SVD can be computed in $\mathcal{O}\!\left(UM\left({D(D+1)}/{2}\right)^{2}\right)\approx\mathcal{O}(UMD^{4})$ FLOP. The total complexity is thus $\mathcal{O}(UMD^{2}+UMD^{4}+D^{3})$, with a worst-case complexity of $\mathcal{O}(UMD^{4})$. On the other hand, for large RIS where $D(D+1)/2>UM$, $\bar{\mathbf{S}}$ is wide, leading to a complexity of $\mathcal{O}\!\left(U^{2}M^{2}{D(D+1)}/{2}\right)\approx\mathcal{O}(U^{2}M^{2}D^{2})$. The total complexity in this regime is $\mathcal{O}(UMD^{2}+U^{2}M^{2}D^{2}+D^{3})$, with a worst-case complexity of $\mathcal{O}(U^{2}M^{2}D^{2})$.
Algorithm 2, for the group-connected RIS, starts with matrix transformations and Kronecker products (lines 1–11), with a combined cost of $\mathcal{O}\!\left(UM\sum_{g=1}^{G}D_{g}^{2}\right)$ complex FLOP. The complexity for the SVD of $\mathbf{J}\in\mathbb{C}^{UM\times\sum_{g=1}^{G}\frac{D_{g}(D_{g}+1)}{2}}$ (line 12) dominates this strategy. When $\mathbf{J}$ is tall, i.e., its number of rows is greater than its number of columns, the SVD requires $\mathcal{O}\!\left(UM\left(\sum_{g=1}^{G}{D_{g}(D_{g}+1)}/{2}\right)^{2}\right)$ FLOP, whereas for a wide matrix $\mathbf{J}$ the complexity becomes $\mathcal{O}\!\left(U^{2}M^{2}\sum_{g=1}^{G}{D_{g}(D_{g}+1)}/{2}\right)$. The final projections (lines 14–18) require $G$ Takagi factorizations, resulting in $\sum_{g=1}^{G}\mathcal{O}(D_{g}^{3})$ FLOP. For equally-sized groups, the worst-case complexities for the tall and wide cases of $\mathbf{J}$ reduce to $\mathcal{O}(UMD^{4}/G^{2})$ and $\mathcal{O}(U^{2}M^{2}D^{2}/G)$, respectively.
The analysis for the random strategy shows that BD-RIS impose a greater computational load compared to a conventional single-connected RIS, which has a linear complexity of $\mathcal{O}(D)$. For the aligned attack with single-connected RIS, iterative optimization methods are often employed [6, 8], which typically have lower per-iteration complexity but may require multiple iterations for convergence. In contrast, Algorithms 1 and 2 are non-iterative, ensuring a feasible runtime. The added cost of BD-RIS represents a trade-off for their superior adversarial performance, as will be seen in Section V.
V Numerical Results
In this section, we evaluate the impact of BD-RIS-induced attacks under both random and optimized strategies and investigate the ability of RSMA to maintain performance under adversarial conditions. We analyze the degradations induced by fully connected, group-connected, and conventional single-connected RIS architectures and provide a comprehensive performance comparison with SDMA.
The simulation setup comprises: (i) a BS equipped with $M=32$ transmit antennas positioned at the origin; (ii) three users located at distances of $d_{1}=30$ m, $d_{2}=50$ m, and $d_{3}=60$ m from the BS, with azimuth angles of $\theta_{1}=25^{\circ}$, $\theta_{2}=15^{\circ}$, and $\theta_{3}=10^{\circ}$; and (iii) an adversarial RIS deployed at a distance of $d_{\text{RIS}}=40$ m from the BS with an azimuth angle of $\theta_{\text{RIS}}=5^{\circ}$. Unless stated otherwise, simulations are performed for $D=200$ reflecting elements, with the group-connected RIS configured as $D_{1}=\cdots=D_{G}=5$, considering a standard imperfect CSI scenario with $\varepsilon^{\text{\tiny BS-U}}=\varepsilon^{\text{\tiny BS-RIS}}=\varepsilon^{\text{\tiny RIS-U}}=0.3$. The small-scale fading for all channel links is modeled following a Rayleigh distribution, representing a rich scattering environment, combined with a power-law path loss model $d^{-\eta}$ with $\eta=3$, where $d$ represents the corresponding distance. In particular, the distance from the RIS to the $u$-th user is obtained by $d_{\text{RIS},u}=\sqrt{d_{\text{RIS}}^{2}+d_{u}^{2}-2d_{\text{RIS}}d_{u}\cos(\theta_{\text{RIS}}-\theta_{u})}$, resulting in $d_{\text{RIS},1}\approx 15.64$ m, $d_{\text{RIS},2}\approx 12.68$ m, and $d_{\text{RIS},3}\approx 20.45$ m. Moreover, the user weights for Algorithms 1 and 2 are set uniformly as $\mu_{1}=\mu_{2}=\mu_{3}=1/3$, the noise power is set to 60-60 dBm, SDMA employs a uniform power allocation with regularized zero-forcing precoders, and RSMA operates under perfect SIC by default. The sensitivity to SIC imperfections, as well as the impact of varying CSI errors, is investigated in Section V-D.
V-A Impact of the RIS Uplink Mode
Figure 3: Impact of the RIS mode in the uplink training phase: (a) average sum rate of RSMA across different RIS architectures; and (b) comparison between RSMA and SDMA under the fully connected RIS. Results are shown for the aligned interference attack.
Fig. 3 investigates the impact of the RIS operation mode during the uplink training phase on the system average sum rate under aligned interference attacks. Fig. 3(a) evaluates this impact across different RIS architectures for the RSMA scheme. It can be seen that using a reflective mode during uplink training significantly boosts the degradation potential for all architectures, but the benefit for the attacker is largest when using BD-RIS. The fully connected RIS, for instance, reduces the sum rate to $23.4$ bits/s/Hz at $40$ dBm, a more severe degradation compared to the $24.7$ bits/s/Hz achieved by the single-connected RIS.
Fig. 3(b) then provides a direct comparison between RSMA and SDMA for the most potent fully connected architecture. The results confirm that the reflective mode is more damaging than the absorption mode for both multiple access schemes. The figure also sheds light on RSMA’s inherent robustness, as it consistently outperforms SDMA under all attacked scenarios. This superior performance is a result of the adaptive power allocation described in Section III-F, which allows RSMA to better manage the channel estimation errors induced by the reflective mode attack. These observations provide the first evidence of RSMA’s robustness advantage over SDMA, which will be further analyzed in the next subsection using the proposed robustness metrics.
V-B Robustness of RSMA
Given that operating in reflective mode during uplink pilot training leads to more severe degradation, we adopt this mode in the subsequent results. The analysis now turns to the robustness of RSMA in comparison to SDMA and across different RIS architectures. Fig. 4(a) presents the robustness index under the fully connected RIS. The results show that RSMA consistently maintains a significantly higher robustness index than SDMA under both random and aligned attacks. For example, under the random interference attack at $40$ dBm, RSMA achieves an index of about $0.88$, whereas SDMA drops to $0.81$. Under the aligned attack, RSMA still recovers to $0.79$ at $40$ dBm, while SDMA reaches only $0.67$. This observation indicates that the gap in robustness between the two schemes widens under the aligned attack, highlighting that RSMA’s performance advantage is even larger when facing severe interference. It is also clear that the aligned attack is consistently more harmful than the random interference, validating the proposed attack strategies.
Complementing the insights above, Fig. 4(b) quantifies the sum-rate degradation of RSMA across different RIS architectures. For all attack types, the degradation curves peak near $12$ dBm and then decrease slightly as transmit power increases, a behavior resulting from RSMA’s flexible interference management mechanism. The figure also confirms that the aligned attack with a fully connected RIS yields the largest degradation among the considered architectures.
Figure 4: Performance degradation and robustness analysis: (a) average robustness index comparing RSMA and SDMA under the fully connected RIS; and (b) average sum-rate degradation of RSMA under random and aligned attacks across different RIS architectures.
V-C Impact of the Number of RIS Reflecting Elements
Fig. 5 analyzes the scaling behavior of the adversarial RIS for a varying number of reflecting elements $D$ under the aligned interference attack, considering a fixed transmit power of $30$ dBm. As shown in Fig. 5(a), the average sum rate for both multiple access schemes degrades as the RIS becomes larger. The performance of SDMA, however, collapses at large $D$, particularly when attacked by a fully connected RIS, whereas the decline of RSMA is considerably less severe. For instance, at $D=400$, SDMA falls below $12$ bits/s/Hz, while RSMA still sustains above $16$ bits/s/Hz when both are attacked by the fully connected RIS.
This trend is also observed in Fig. 5(b) in terms of sum-rate degradation, where the gap in degradation between RSMA and SDMA clearly widens as $D$ increases. These results highlight the scaling advantage of the different RIS architectures for an attacker, as the performance curves for the three architectures become more separated with larger $D$, demonstrating that BD-RIS become increasingly advantageous over a conventional RIS. At the same time, the results confirm the scalability of RSMA in adversarial environments, since it continues to offer performance gains over SDMA even when the adversarial RIS dimension increases. Next, we examine how RSMA’s robustness is affected by practical impairments.
Figure 5: Impact of the number of reflecting elements $D$ for different RIS architectures: (a) average sum rate; and (b) average sum-rate degradation. Results assume a fixed transmit power of $30$ dBm under the aligned interference attack.
V-D Impact of CSI and SIC Imperfections
Figure 6: Impact of imperfect CSI on the system sum rate: (a) versus channel error $\epsilon^{\text{\tiny BS-U}}$ in the BS-user link, with $\epsilon^{\text{\tiny BS-RIS}}=\epsilon^{\text{\tiny RIS-U}}=0.3$ for the RIS-associated channels at the attacker; and (b) versus $\epsilon^{\text{\tiny BS-RIS}}=\epsilon^{\text{\tiny RIS-U}}$ at the attacker, with a fixed error of $\epsilon^{\text{\tiny BS-U}}=0.3$ in the BS-user link. Results are shown for both random and aligned attacks induced by the fully connected RIS, with a transmit power of $30$ dBm.
Fig. 6 provides a deeper analysis of the impact of imperfect CSI, considering a fixed transmit power of $30$ dBm for both random and aligned attacks induced by the fully connected RIS. Fig. 6(a) investigates system performance versus the error in the legitimate BS-user link $\epsilon^{\text{\tiny BS-U}}$, while the attacker’s CSI for the RIS-related links has a fixed error of $\epsilon^{\text{\tiny BS-RIS}}=\epsilon^{\text{\tiny RIS-U}}=0.3$. As can be seen, at $\epsilon^{\text{\tiny RIS-U}}=0$, RSMA and SDMA exhibit identical performance under the same attack. This behavior is explained by the fact that RSMA reduces to SDMA when the BS has perfect channel knowledge. As the error in the legitimate channel increases, the sum rate of SDMA under both attack types monotonically decreases. RSMA, in contrast, shows a slight initial increase in its sum rate for small $\epsilon^{\text{\tiny BS-U}}>0$ under both random and aligned attacks.
Fig. 6(b) examines the opposite scenario, where the legitimate BS-user CSI error is fixed at $\epsilon^{\text{\tiny BS-U}}=0.3$, and the attacker’s CSI for the RIS-related links is varied. The results show that as the attacker’s channel estimation error increases, the performance of both RSMA and SDMA under attack improves. This is because higher estimation errors at the adversary lead to a less effective optimization for the aligned attack and a less coherent random channel for the random attack, resulting in less damaging interference. Across all scenarios where the legitimate channel is imperfect, i.e., $\epsilon^{\text{\tiny BS-U}}>0$, RSMA maintains a clear sum rate advantage over SDMA.
Figure 7: Impact of imperfect SIC on the system sum rate: (a) versus transmit power for different SIC error factors $\xi$; and (b) versus the SIC error factor $\xi$ for a fixed transmit power of $30$ dBm. Results assume the aligned interference attack induced by the fully connected RIS.
Finally, Fig. 7 investigates the system’s sensitivity to SIC imperfections under the aligned interference attack. Fig. 7(a) illustrates the average sum rate versus transmit power for different SIC error factors $\xi$. As expected, the performance of SDMA is unaffected by this parameter. The results for RSMA, however, reveal a high sensitivity to SIC quality. Even a small error factor of $\xi=10^{-4}$ causes a noticeable degradation compared to the perfect SIC case ($\xi=0$). For a larger error of $\xi=10^{-3}$, the performance of RSMA collapses, with its sum rate saturating at high power levels and nearly coinciding with that of SDMA. This indicates that RSMA’s ability to recover from the attack is highly dependent on the quality of the SIC implementation.
The impact of the SIC error factor is further detailed in Fig. 7(b) for a fixed transmit power of $30$ dBm. The performance of RSMA shows a steep decline as $\xi$ increases, for both the safe and attacked scenarios. In particular, RSMA approaches the performance of SDMA from $\xi=0.6\times 10^{-3}$ onward. On one hand, this exposes the sensitivity of RSMA to imperfect SIC. On the other hand, it reveals the effectiveness of its adaptive power allocation, which, being aware of SIC errors, reallocates more power to private messages so that RSMA at least matches SDMA under imperfect SIC. This finding highlights a critical trade-off: while RSMA is fundamentally more robust to external CSI uncertainties, its overall performance advantage is contingent on the availability of high-quality hardware for interference cancellation.
VI Conclusions
This paper has investigated the robustness of RSMA-based multi-user MISO systems to adversarial attacks induced by BD-RIS. Two effective attack strategies were described and studied: random and aligned interference. For each, novel algorithms were proposed to generate valid reflection coefficients respecting the architectural BD-RIS constraints. Extensive simulations revealed that RSMA demonstrated significantly greater robustness to these attacks than the benchmark SDMA scheme, especially under imperfect CSI. Moreover, BD-RIS significantly enhances adversarial efficacy compared to conventional RIS, with particularly severe performance degradation observed when the BD-RIS manipulates the uplink training phase via reflective operation. Future work should expand this analysis and investigate the adversarial potential of multi-sector BD-RIS and its non-reciprocal architectures. The development of robust countermeasures against such threats is another crucial direction.
Acknowledgments
This work was supported by the Commonwealth Cyber Initiative (cyberinitiative.org) in Virginia, US, an investment to advance cyber R&D, innovation, and workforce development, the National Science Foundation under grants no. 2326599 and 231879, and the Research Council of Finland through the 6G-ConCoRSe project (grant no. 359850) and the 6G Flagship programme (grant no. 369116).